# Kang Zhu1,2, Qianbo Zang911footnotemark: 1, Shian Jia411footnotemark: 1, Siwei Wu1,311footnotemark: 1, Feiteng Fang6,1011footnotemark: 1 Yizhi Li1,3,Shuyue Guo1, Tianyu Zheng1,2, Bo Li5, Haoning Wu2, Xingwei Qu1

### v1

@@ Abstract
With the remarkable success achieved by Multimodal Large Language Models (MLLMs), numerous benchmarks have been designed to assess MLLMs’ ability to guide their development in image perception tasks (e.g., image captioning and visual question answering). However, the existence of numerous benchmarks results in a substantial computational burden when evaluating model performance across all of them. Moreover, these benchmarks contain many overly simple problems or challenging samples, which do not effectively differentiate the capabilities among various MLLMs. To address these challenges, we propose a pipeline to process the existing benchmarks, which consists of two modules (i.e., Semi-Automated Screening Process and Eliminating Answer Leakage). The Semi-Automated Screening Process filters out samples that cannot distinguish the model’s capabilities by synthesizing various MLLMs and manually evaluating them. The Eliminate Answer Leakage filters the samples whose answers can be inferred without images. Finally, we curate the LIME-M: Less Is More for Evaluation of Multimodal LLMs, which is a lightweight Multimodal benchmark and can more effectively evaluate the performance of different models. Our experiments demonstrate that 1. LIME-M can better distinguish the performance of different MLLMs within less sample numbers (24% of original) and time spent (23% of original); 2. LIME-M eliminates answer leakage and mainly focuses on the information within images; 3. The current automatic metric (i.e., CIDEr) is insufficient for evaluating MLLMs’ capabilities in captioning. Moreover, we find that removing the caption task score when calculating the overall score demonstrates a more precise reflection of model performance differences. All our codes and data are released at https://github.com/kangreen0210/LIME-M
@@ 1 Introduction
However, existing MLLM benchmarks and unified evaluation frameworks cannot effectively and efficiently reflect the ability of MLLMs. Current benchmarks include numerous relatively simple samples (i.e., how many chairs are in the picture) and some incorrect questions caused by annotation issues. Most MLLMs will consistently perform on these samples (i.e., all correct or all wrong). Therefore, those benchmarks cannot fully reflect the gap between different MLLMs and across various tasks. Besides, the current unified Multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks. How to select effective evaluation data has been largely overlooked by current researchers.
To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME-M, which contains 9403 samples and is refined across 10 tasks within 6 domains. We select six major tasks in the Multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain. To eliminate bias introduced by individual models, we chose 9 models as judges and filter samples based on their performance. On the one hand, we removed samples that most models answered correctly due to the fact that they cannot distinguish the capabilities among different models. On the other hand, we use a method that combines humans and MLLMs to filter out some abnormally difficult samples. Meanwhile, we use LLMs to filter out samples that could be answered directly from the context of the question. After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME-M).
We conducted various experiments on LIME-M using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment. We have made several valuable findings:
  • LIME-M can better reflect the performance differences of MLLMs. On our LIME-M benchmark, under consistent conditions (same model series, same model size), different MLLMs demonstrate a wider score range, indicating that LIME-M is more effective at reflecting performance differences between models with a smaller amount of data.
@@ 2 Method
Figure 2 shows the statistics of different subtasks within our LIME-M benchmark. It is worth mentioning that the proportion of easy and wrong samples exceeds 30%. Out of the 10 subtasks, 6 have proportions exceeding 50%. Notably, for the POPE dataset, 95% of the data can be classified as noisy or erroneous. This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
To address those issues, we propose a general data curation pipeline, which consists of three main stages: (1) Open-source models as judges, (2) Semi-automated screening process, and (3) Eliminating Answer Leakage. Our approach aims to enhance existing benchmarks by removing inaccurate and oversimplified data.
@@ 2.3 Eliminating Answer Leakage
3 LIME-M: A comprehensive MLLMs benchmark
@@
Table 1: Data statics: Full Size: the size of the original dataset, Lite Size: the final size of the LIME-V. For the COCO-Caption dataset, we selected the 2017 subset, and for the ScienceQA dataset, we chose the ScienceQA-IMG subset.
In this section, we propose LIME-M, a comprehensive benchmark for Multimodal Large Language Models (MLLMs). LIME-M streamlines existing mainstream benchmarks. Table 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning. For each sub-dataset, we aim to keep the size around 1k samples.
@@ 3.1 Task definition
LIME-M is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000. After our three-stage data curation, the data scale of each benchmark is significantly reduced. Figure 1 shows the number of samples removed at each stage compared to the original dataset. The amount of data removed varies at each stage, with the most being removed in the first stage, reflecting a large number of low-difficulty or data-leakage samples in the existing 9 MLLMs. Comparing the data volumes before and after the second stage of semi-automated screening, we can see that many datasets, such as OK-VQA and TextVQA, have a high rate of low-quality data leading to MLLMs’ incorrect answers. Additionally, some datasets, such as ScienceQA and AI2D, have a significant amount of data removed after the third stage, indicating that many questions in these datasets can be answered correctly by MLLMs without relying on images. The statistics of the curated data are shown in  1. LIME-M includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
@@ 4.1 Experiment Setting
To evaluate the quality of LIME-M, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
  • Main experiment: To demonstrate the performance of LIME-M, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
@@
  • Text-only question with Image Description set: Image Description (ID) refers to simple descriptions of images that represent superficial information contained within them. For most MLLMs, questions containing only superficial information are easy to answer; inversely, questions requiring complex visual inference are significantly more challenging. To further validate whether LIME-M can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
@@ 4.2 Baselines
We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2  [^1] https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline. To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME-M. For models not supported by this framework, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
@@
For most tasks included in LIME-M, we reference the metrics computation methods used in MLLMs-eval. Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses. For tasks such as OK-VQA and TextVQA, the metric scores are calculated based on the overlap between the response and the candidate answers. For tasks like TextCaps and COCO-Caption2017, we use the CIDEr as the score. The ANLS metric is used for the infoVQA task, and the Relaxed Overall metric is used for the ChartQA task.
@@ 5.1 Main Result
T/F Common VQA InfoVQA ScienceQA OCR Captioning
Model Size Overall POPE $\uparrow$ TVQA $\uparrow$ OK VQA $\uparrow$ ChQA $\uparrow$ IVQA $\uparrow$ AI2D $\uparrow$ SciQA $\uparrow$ OCRBen $\uparrow$ COCO $\uparrow$ TCaps $\uparrow$
InternVL-2 2023 ( - ) 40B 64.19 51.69 77.98 19.45 88.33 73.92 69.20 98.33 34.90 63.10 48.94
Qwen2-VL 2023b ($\uparrow$ 1) 7B 62.56 53.05 74.56 28.18 83.17 78.50 58.20 75.00 35.70 68.74 110.60
InternVL-1.5 2024 ($\downarrow$ 1) 26B 62.00 51.69 69.88 35.47 87.00 70.92 54.81 91.67 27.90 69.24 112.63
InternVL-2 2023 ( - ) 26B 61.48 54.63 75.20 16.08 87.67 70.58 62.80 96.33 32.70 76.18 56.91
LLaVA-OneVision 2024 ( - ) 7B 60.28 52.37 65.22 37.32 80.83 68.17 59.20 92.00 22.00 104.74 108.18
InternVL-2 2023 ( - ) 8B 59.52 49.21 66.10 24.20 83.75 68.25 60.40 95.67 32.50 42.55 25.44
XComposer2-4KHD 2023 ($\uparrow$ 4) 7B 55.67 46.28 60.30 28.13 80.42 66.17 54.90 70.00 24.40 97.07 78.07
InternVL-2 2023($\downarrow$ 1) 4B 54.86 47.18 62.29 17.48 81.92 60.50 54.00 94.33 30.90 35.99 21.67
CogVLM-2 2024 ($\uparrow$ 6) 19B 52.02 51.02 69.46 4.92 80.33 59.50 45.00 84.00 31.40 23.67 34.01
Qwen2-VL 2023b ($\uparrow$ 5) 2B 51.59 50.79 70.70 16.87 67.50 65.00 42.90 58.00 31.30 75.06 101.72
InternVL-2 2023 ($\downarrow$ 2) 2B 51.25 50.79 59.56 21.87 71.75 54.00 45.80 92.33 31.00 51.95 42.59
CogVLM-1 2023 ($\uparrow$ 1) 17B 49.53 55.10 71.20 31.70 61.67 57.25 31.40 53.67 19.30 29.28 38.56
Cambrian 2024 ($\downarrow$ 5) 34B 48.78 49.44 57.28 22.03 71.83 43.17 54.80 78.33 18.00 4.27 4.97
Cambrian 2024 ($\downarrow$ 4) 13B 47.06 50.79 58.93 24.13 69.25 42.83 45.20 62.67 19.50 7.40 6.52
Cambrian 2024 ($\downarrow$ 3) 8B 46.41 49.89 59.00 25.23 69.42 37.67 43.60 69.00 19.80 5.85 6.41
InternVL-2 2023 ($\uparrow$ 2) 1B 45.88 52.82 57.55 15.37 65.83 46.25 37.00 84.33 29.90 15.19 13.19
LLaVA-1.6 2024 ($\uparrow$ 3) 34B 42.73 47.00 51.20 10.40 64.33 42.08 49.60 64.67 17.10 84.25 48.25
MiniCPM-LLaMA3-2.5 2024 ($\downarrow$ 3) 8B 40.62 43.10 61.80 25.30 69.00 48.10 6.60 6.60 25.70 31.91 39.86
LLaVA-OneVision 2024 ($\uparrow$ 4) 0.5B 39.85 48.98 48.61 23.13 55.42 40.67 31.20 53.00 19.70 96.40 90.28
LLaVA-LLaMA3 2023 ($\downarrow$ 3) 8B 39.81 44.24 40.01 34.72 64.42 22.25 40.72 61.67 13.90 99.35 48.71
Mantis-Idefics-2 2024 ( - ) 8B 38.09 44.24 44.51 29.07 59.00 20.50 35.80 70.00 14.80 61.82 103.07
Deepseek-VL 2024 ($\uparrow$ 2) 7B 37.17 48.50 44.80 25.00 54.67 22.33 37.00 68.33 11.70 54.22 83.21
LLaVA-1.6-vicuna 2024 ($\downarrow$ 4) 13B 35.93 43.10 43.90 16.10 54.50 28.75 38.10 53.00 14.70 76.62 47.83
Idefics-2 2024 ($\downarrow$ 2) 8B 34.85 42.00 56.50 35.60 13.08 23.92 38.10 78.67 19.60 61.23 94.51
LLaVA-1.6-vicuna 2024 ( - ) 7B 29.27 41.10 39.00 12.50 43.08 22.67 27.10 47.33 10.90 76.05 48.35
Mantis-SigLIP 2024 ($\uparrow$ 1) 8B 28.70 45.60 26.34 32.43 35.33 16.25 27.70 62.67 4.90 68.16 81.21
MiniCPM 2024 ($\uparrow$ 2) 1.0 25.58 44.00 37.00 6.20 35.75 13.42 27.90 60.67 6.70 68.96 76.65
LLaVA-1.5 2023a ($\downarrow$ 2) 13B 20.12 36.60 19.50 32.10 5.50 12.42 25.90 48.33 2.70 80.89 68.73
LLaVA-1.5 2023a ($\downarrow$ 1) 7B 16.98 32.51 16.50 23.43 5.25 9.08 24.05 49.00 2.20 79.20 65.73
InstructBLIP-vicuna 2023 ( - ) 7B 15.30 45.10 11.40 22.10 3.00 9.08 21.90 34.33 2.00 102.08 53.14
Tiny-LLaVA-1 2024 ( - ) 1.4B 13.76 37.00 18.70 0.90 4.50 12.17 22.80 44.67 1.60 63.19 58.91
InternVL-2 40B 80.31 89.23 82.59 50.98 85.52 76.08 85.88 98.56 79.90 99.15 62.03
InternVL-1.5 26B 79.49 88.90 79.00 60.70 83.70 72.50 78.90 94.50 71.40 95.80 148.10
Qwen2-VL 7B 79.14 88.17 80.92 55.68 83.32 78.86 80.73 85.57 81.20 92.13 144.36
InternVL-2 26B 78.82 88.64 82.06 48.50 84.44 72.72 83.16 97.47 77.60 110.30 80.10
LLaVA-OneVision 7B 78.71 89.17 76.02 60.98 80.12 70.69 81.38 95.88 62.10 140.45 136.97
InternVL-2 8B 77.84 87.90 77.00 52.02 82.48 70.65 82.25 97.03 76.50 89.77 36.70
InternVL-2 4B 73.97 87.71 74.51 38.43 81.04 65.19 78.08 96.03 75.00 54.08 30.17
Cambrian 34B 73.26 88.46 72.11 52.07 74.60 51.48 80.41 85.52 59.00 8.18 6.08
InternVL-2 2B 73.00 88.90 72.39 43.74 74.72 57.69 72.70 94.25 75.50 79.52 59.81
Cambrian 13B 72.39 88.53 73.07 53.28 72.60 50.73 73.93 79.08 61.40 14.33 9.44
XComposer-4KHD 7B 71.93 87.00 74.30 51.90 80.60 72.80 34.40 96.00 66.90 134.00 111.40
Cambrian 8B 71.84 88.24 72.47 52.17 73.44 48.05 72.99 80.32 61.60 9.13 7.97
CogVLM-1 17B 71.34 88.90 79.70 46.90 67.00 63.30 61.90 70.50 59.10 28.40 44.70
MiniCPM-LLaMA3-2.5 8B 71.22 88.00 75.00 52.30 72.90 56.90 71.70 53.00 69.40 35.50 52.90
Qwen2-VL 2B 70.86 87.78 78.70 40.59 73.12 67.15 70.21 77.89 75.30 103.52 131.92
CogVLM-2 19B 69.93 87.56 77.59 18.51 79.84 62.62 72.41 90.93 76.60 24.10 42.23
LLaVA-LLaMA3 8B 69.74 87.80 65.40 60.20 69.30 37.60 71.60 73.30 55.00 135.00 69.60
InternVL-2 1B 68.46 87.94 69.67 33.84 71.40 52.02 62.56 89.59 74.20 49.34 18.03
LLaVA-1.6-vicuna 13B 67.29 87.50 67.00 46.30 62.20 41.50 70.40 73.50 55.00 101.90 67.30
LLaVA-1.6 34B 67.22 85.60 68.90 31.00 67.40 51.90 76.10 82.70 58.60 114.40 69.10
Mantis-Idefics-2 8B 66.91 86.90 63.51 52.50 63.56 31.17 66.81 81.80 54.20 79.42 134.08
Idefics-2 8B 66.73 86.80 71.30 53.90 26.40 37.00 69.20 87.20 61.60 71.90 119.10
LLaVA-OneVision 0.5B 65.65 88.33 65.85 44.17 61.36 46.23 57.09 67.03 57.60 131.90 120.81
Deepseek-VL 7B 65.62 87.10 63.20 48.70 60.60 34.30 63.40 81.70 43.30 67.60 110.10
LLaVA-1.6-vicuna 7B 64.80 87.60 64.90 44.20 55.00 37.00 65.30 70.20 52.40 100.00 72.00
LLaVA-1.5 13B 59.58 87.10 48.70 58.30 18.10 29.50 59.40 72.80 33.60 115.40 104.00
Mantis-SigLIP 8B 58.96 81.47 49.59 52.90 42.56 26.56 57.84 75.36 34.50 91.37 111.43
LLaVA-1.5 7B 57.27 87.00 46.10 53.40 18.20 25.80 55.20 69.50 31.50 109.00 98.00
MiniCPM 1.0 56.18 85.10 55.30 47.30 15.40 20.10 56.90 43.00 60.00 25.90 41.60
InstructBLIP-vicuna 7B 47.87 85.00 33.20 45.20 12.50 22.90 34.00 36.40 25.90 141.40 74.00
Tiny-LLaVA-1 1.4B 34.30 56.30 38.50 3.80 11.10 22.20 32.30 58.20 17.20 80.90 83.10
Table 2: Comparing overall scores of LIME-M and Original. Top: results on LIME-M, Bottom: results on the original dataset. The arrow next to the model name indicates the change in ranking on LIME-M compared to the original dataset. $\uparrow$: upward shift, $\downarrow$: downward shift, and -: no change.
As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME-M benchmark. Overall, models with larger parameter sizes and newer versions tend to have higher overall scores. The InternVL-1.5, InternVL-2-Large (26B,40B) and LLaVA-OneVision-7B achieve the best overall performance, whose overall scores all surpass 60%. The performance of InternVL-2-Small(1B-8B), CogVLM series, and Cambrian series are second only to InternVL-2-Large and LLaVA-OneVision-7B, with their overall scores ranging from 45% to 60%. Comparing the overall score of LIME-M and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline. Conversely, the CogVLM and LLaVA-OneVision series’ overall scores show an improvement. Specifically, that of CogVLM2 and XComposer-4KHD experience significant fluctuations, improved by 4% and 6% respectively
@@
Figure 4: Correlation distribution between LIME-M and Wildvison Elo.

### v2

# King Zhu1,2, Qianbo Zang911footnotemark: 1, Shian Jia411footnotemark: 1, Siwei Wu1,311footnotemark: 1, Feiteng Fang6,811footnotemark: 1, Yizhi Li1,3 Shawn Gavin1, Tuney Zheng1,2, Jiawei Guo1,2, Bo Li5, Haoning Wu2, Xingwei Qu1

@@ Abstract
Multimodal Large Language Models (MLLMs) are measured on numerous benchmarks like image captioning, visual question answering, and reasoning. However, these benchmarks often include overly simple or uninformative samples, making it difficult to effectively distinguish the performance of different MLLMs. Additionally, evaluating models across many benchmarks creates a significant computational burden. To address these issues, we propose LIME (Less Is More for MLLM Evaluation), a refined and efficient benchmark curated using a semi-automated pipeline. This pipeline filters out uninformative samples and eliminates answer leakage by focusing on tasks that require image-based understanding. Our experiments show that LIME reduces the number of samples by 76% and evaluation time by 77%, while more effectively distinguishing between models. Notably, we find that traditional automatic metrics like CIDEr are insufficient for evaluating MLLMs’ captioning performance, and excluding the caption task score yields a more accurate reflection of overall model performance. All code and data are available at https://github.com/kangreen0210/LIME.
@@ 1 Introduction
However, existing MLLM benchmarks and unified evaluation frameworks cannot effectively and efficiently reflect the ability of MLLMs. Current benchmarks include numerous relatively simple samples (i.e., how many chairs are in the picture) and some incorrect questions caused by annotation issues. Most MLLMs will consistently perform on these samples (i.e., all correct or all wrong). Therefore, those benchmarks cannot fully reflect the gap between different MLLMs and across various tasks. Besides, the current unified Multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks. The selection of effective evaluation data has been largely overlooked by current researchers.
To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME, which contains 9403 samples and is refined across 10 tasks within 6 domains. We select six major tasks in the Multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain. To eliminate bias introduced by individual models, we chose 9 models as judges and filter samples based on their performance. On the one hand, we removed samples that most models answered correctly due to the fact that they cannot distinguish the capabilities among different models. On the other hand, we use a method that combines humans and MLLMs to filter out some abnormally difficult samples. Meanwhile, we use LLMs to filter out samples that could be answered directly from the context of the question. After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME).
We conducted various experiments on LIME using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment. We have made several valuable findings:
  • LIME can better reflect the performance differences of MLLMs. On our LIME benchmark, under consistent conditions (same model series, same model size), different MLLMs demonstrate a wider score range, indicating that LIME is more effective at reflecting performance differences between models with a smaller amount of data.
@@ 2 Method
Figure 2 shows the statistics of different subtasks within our LIME benchmark. It is worth mentioning that the proportion of easy and wrong samples exceeds 30%. Out of the 10 subtasks, 6 have proportions exceeding 50%. Notably, for the POPE dataset, 95% of the data can be classified as noisy or erroneous. This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
Inspired by MMStar (Chen et al., 2024a), we utilize open-source MLLMs and LLMs as the judges for filtering, specifically, we removing the existing annotation errors. The overall pipeline consists of three main stages: (1) Using open-source models as judges, (2) A semi-automated screening process, and (3) Eliminating answer leakage. Our approach aims to improve existing benchmarks by removing inaccurate and oversimplified data.
@@ 2.3 Eliminating Answer Leakage
3 LIME: A comprehensive MLLMs benchmark
@@
Table 1: Data statics: Full Size: the size of the original dataset, Lite Size: the final size of the LIME. For the COCO-Caption dataset, we selected the 2017 subset, and for the ScienceQA dataset, we chose the ScienceQA-IMG subset.
In this section, we propose LIME, a comprehensive benchmark for Multimodal Large Language Models (MLLMs). LIME streamlines existing mainstream benchmarks. Table 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning. For each sub-dataset, we aim to keep the size around 1k samples.
@@ 3.1 Task definition
LIME is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000. After our three-stage data curation, the data scale of each benchmark is significantly reduced. Figure 1 shows the number of samples removed at each stage compared to the original dataset. The amount of data removed varies at each stage, with the most being removed in the first stage, reflecting a large number of low-difficulty or data-leakage samples in the existing 9 MLLMs. Comparing the data volumes before and after the second stage of semi-automated screening, we can see that many datasets, such as OK-VQA and TextVQA, have a high rate of low-quality data leading to MLLMs’ incorrect answers. Additionally, some datasets, such as ScienceQA and AI2D, have a significant amount of data removed after the third stage, indicating that many questions in these datasets can be answered correctly by MLLMs without relying on images. The statistics of the curated data are shown in  1. LIME includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
@@ 4.1 Experiment Setting
To evaluate the quality of LIME, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
  • Main experiment: To demonstrate the performance of LIME, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
@@
  • Text-only question with Image Description set: Image Description (ID) refers to simple descriptions of images that represent superficial information contained within them. For most MLLMs, questions containing only superficial information are easy to answer; inversely, questions requiring complex visual inference are significantly more challenging. To further validate whether LIME can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
@@ 4.2 Baselines
We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2  [^1] https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024b) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline. To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME. For models not supported by this framework, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
@@
For most tasks included in LIME, we reference the metrics computation methods used in MLLMs-eval. Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses. For tasks such as OK-VQA and TextVQA, the metric scores are calculated based on the overlap between the response and the candidate answers. For tasks like TextCaps and COCO-Caption2017, we use the CIDEr as the score. The ANLS metric is used for the infoVQA task, and the Relaxed Overall metric is used for the ChartQA task.
@@ 5.1 Main Result
Model Size LIME Original Reasoning VQA InfoQA SciQA OCR Caption
InternVL-2 2023 40B 64.15 ( - ) 80.31 51.69 48.72 81.12 75.92 34.90 56.02
Qwen2-VL 2023b 7B 62.53 ($\uparrow$ 1) 79.14 53.05 51.37 80.83 62.08 35.70 89.67
InternVL-1.5 2024b 26B 61.97 ($\downarrow$ 1) 79.49 51.69 52.68 78.96 63.32 27.90 90.93
InternVL-2 2023 26B 61.46 ( - ) 78.82 54.63 45.64 79.12 70.54 32.70 66.54
LLaVA-OneVision 2024 7B 60.25 ( - ) 78.71 52.37 51.27 74.50 66.77 22.00 106.46
InternVL-2 2023 8B 59.49 ( - ) 77.84 49.21 45.15 76.00 68.54 32.50 34.00
XComposer2-4KHD 2023 7B 55.64 ($\uparrow$ 4) 71.93 46.28 44.22 73.29 58.38 24.40 87.57
InternVL-2 2023 4B 54.84 ($\downarrow$ 1) 73.97 47.18 39.89 71.21 63.31 30.90 28.83
CogVLM-2 2024 19B 52.02 ($\uparrow$ 6) 69.93 51.02 37.19 69.92 54.00 31.40 28.84
Qwen2-VL 2023b 2B 51.59 ($\uparrow$ 5) 70.86 50.79 43.78 66.25 46.38 31.30 88.39
InternVL-2 2023 2B 51.25 ($\downarrow$ 2) 73.00 50.79 40.71 62.88 56.54 31.00 47.27
CogVLM-1 2023 17B 49.55 ($\uparrow$ 1) 71.34 55.10 51.45 59.46 36.54 19.30 33.92
Cambrian 2024 34B 48.79 ($\downarrow$ 5) 73.26 49.44 39.66 57.50 60.23 18.00 4.62
Cambrian 2024 13B 47.07 ($\downarrow$ 4) 72.39 50.79 41.53 56.04 49.23 19.50 6.96
Cambrian 2024 8B 46.42 ($\downarrow$ 3) 71.84 49.89 42.12 53.55 49.46 19.80 6.13
InternVL-2 2023 1B 45.90 ($\uparrow$ 2) 68.46 52.82 36.46 56.04 47.92 29.90 14.19
LLaVA-1.6 2024 34B 42.74 ($\uparrow$ 3) 67.22 47.00 30.80 53.21 53.08 17.10 66.25
MiniCPM-LLaMA3-2.5 2024 8B 40.63 ($\downarrow$ 3) 71.22 43.10 43.55 58.55 6.60 25.70 35.89
LLaVA-OneVision 2024 0.5B 39.88 ($\uparrow$ 4) 65.65 48.98 35.87 48.04 36.23 19.70 93.34
LLaVA-LLaMA3 2023 8B 39.82 ($\downarrow$ 3) 69.74 44.24 37.36 43.33 45.56 13.90 74.03
Mantis-Idefics-2 2024 8B 38.11 ( - ) 66.91 44.24 36.79 39.75 43.69 14.80 82.44
Deepseek-VL 2024 7B 37.20 ($\uparrow$ 2) 65.62 48.50 34.90 38.50 44.23 11.70 68.72
LLaVA-1.6-vicuna 2024 13B 35.95 ($\downarrow$ 4) 67.29 43.10 30.00 41.63 41.54 14.70 62.23
Idefics-2 2024 8B 34.88 ($\downarrow$ 2) 66.73 42.00 46.05 18.50 47.46 19.60 77.87
LLaVA-1.6-vicuna 2024 7B 29.30 ( - ) 64.80 41.10 25.75 32.88 31.77 10.90 62.20
Mantis-SigLIP 2024 8B 28.76 ($\uparrow$ 1) 58.96 45.60 29.39 25.79 35.77 4.90 74.69
MiniCPM 2024 1.0 25.63 ($\uparrow$ 2) 56.18 44.00 21.60 24.58 35.46 6.70 72.80
LLaVA-1.5 2023a 13B 20.17 ($\downarrow$ 2) 59.58 36.60 25.80 8.96 31.08 2.70 74.81
LLaVA-1.5 2023a 7B 17.03 ($\downarrow$ 1) 57.27 32.51 19.97 7.17 29.81 2.20 72.47
InstructBLIP-vicuna 2023 7B 15.39 ( - ) 47.87 45.10 16.75 6.04 24.77 2.00 77.61
Tiny-LLaVA-1 2024 1.4B 13.83 ( - ) 34.30 37.00 9.80 8.33 27.85 1.60 61.05
Table 2: Comparing overall scores of LIME and Original. The arrow next to the LIME score indicates the change in ranking on LIME compared to the original dataset. $\uparrow$: upward shift, $\downarrow$: downward shift, and -: no change.
As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME benchmark. Overall, models with larger parameter sizes and newer versions tend to have higher overall scores. The InternVL-1.5, InternVL-2-Large (26B,40B) and LLaVA-OneVision-7B achieve the best overall performance, whose overall scores all surpass 60%. The performance of InternVL-2-Small(1B-8B), CogVLM series, and Cambrian series are second only to InternVL-2-Large and LLaVA-OneVision-7B, with their overall scores ranging from 45% to 60%. Comparing the overall score of LIME and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline. Conversely, the CogVLM and LLaVA-OneVision series’ overall scores show an improvement. Specifically, that of CogVLM2 and XComposer-4KHD experience significant fluctuations, improved by 4% and 6% respectively
@@
Figure 4: Correlation distribution between LIME and Wildvison Elo.